A laser emits light with a frequency of 1.59 Times
Answer #1

E = h \times \nu

h = Planck's constant = 6.626 \times 10^{-34} Js

\nu = frequency = 1.59 \times 10^{15} Hz

E = 6.626 \times 10^{-34} \times 1.59 \times 10^{15} J = 1.053 \times 10^{-18} J.